### v1

# LLM Agents for Automated Dependency Upgrades

Abstract
As a codebase expands over time, its library dependencies can become outdated and require updates to maintain innovation and security. However, updating a library can introduce breaking changes in the code, necessitating significant developer time for maintenance. To address this, we introduce a framework of LLM agents to be used in combination with migration documentation to automatically recommend and apply code updates and ensure compatibility with new versions. Our solution can automatically localize updated library usages in live Java codebases and implement recommended fixes in a user-friendly manner. The system architecture consists of multiple key components: a Summary Agent, Control Agent, and Code Agent. To validate our approach, we apply the framework on an industrial use case by which we create three synthetic code repositories with major Upgrade changes and benchmark our approach against state-of-the-art methods. Results show that our approach not only performs upgrades using fewer tokens across all cases but also achieves a precision of 71.4%, highlighting its efficiency and effectiveness compared to state-of-the-art methods.
I Introduction
Modern software projects strive to avoid reinventing the wheel by leveraging reusable functionality provided by open-source libraries. However, updating these libraries can introduce breaking changes in the code, posing challenges for developers [1]. Many software projects heavily depend on open-source libraries that offer ready-to-use functionality, including those available for Java. For example, Maven Central is a repository for Java libraries, providing developers with access to a vast array of libraries for their projects [2]. While these libraries offer convenience and accessibility, they also introduce dependencies, and frequent library updates can result in breaks in developers’ code.
In the Java ecosystem, contributors often use warnings to alert library users of upcoming changes to the API and functionality. These warnings are typically issued several releases before the actual changes take effect, allowing developers time to update their code for compatibility [3]. However, there is currently no clearly established method for automatically updating library usages in code to account for changes in Java libraries, leading to costly and time-consuming manual maintenance without immediate incentives. This issue is particularly pronounced in larger codebases, where developers might ignore warnings or lock into a specific library version [4, 5]. Such strategies can hinder a team’s ability to innovate in the long term, preventing them from benefiting from new features, performance improvements, and bug fixes offered by updated libraries.
Some automated solutions have been proposed in the past, but have not been widely adopted, giving way to further research opportunities in this domain. Furthermore, large language models (LLMs) have been at the forefront of AI, and the software engineering domain is no exception. LLM systems have been used to tackle a number of coding challenges, with a focus on code generation and code completion [6], or automated program repair [7]. Nevertheless, the maintenance and evolution of code remains a critical task in engineering, impacting code quality and reusability [8]. As such, in this paper we propose a framework of LLM Agents for automated dependency upgrades. This system builds upon our previous work, the Autonomous LLM-based Multi-Agent Software Engineering Framework (ALMAS) [9]. The system is evaluated on an industrial use case with the aim to update code repositories after completing a major library upgrade.
Figure 1: Diagram of the multi-agent LLM framework for automated Java dependency upgrades, showing the process of code summarization, change localization, and implementation, with iterative compilation and testing, and handover to a human.
II Related Work
The notion of automatically updating a codebase to ensure compatibility with newer dependencies has been explored previously, with a focus on identifying differences between library versions. Dagenais and Robillard introduced SemDiff [10], a tool that recommends adaptations to client programs by examining how a framework adjusted to its own modifications. The concept is that if a method is removed from a library, one can pinpoint all callers of that method and assess how they adapted to its removal, thereby proposing a similar solution for the developer’s codebase. Similarly, Nguyen et al. developed LIBSYNC, a tool that assists developers in adapting API usage code by learning complex API usage adaptation patterns from other clients that have already transitioned to a new library version [11].
A key aspect of most of these studies is their focus on statically typed programming languages. However, more recent research has addressed dynamically typed languages. For example, Nielsen et al. aimed to identify locations affected by breaking changes in a codebase using a collection of semantic patches in JavaScript [12]. Various other methodologies for processing code and related documentation have been explored, including manual or rule-based labelling of changelogs, basic NLP techniques like TF-IDF and word embeddings for matching error messages with relevant documentation [13], and popular end-to-end neural approaches.
One dominant approach was neural machine translation (NMT), adapted from NLP, which frames code-related tasks such as bug fixing [14], code summarization [15], or code review automation [16] as sequence-to-sequence prediction tasks using deep learning architectures like encoder-decoder LSTM [17] [18] or Transformer [19]. Despite the versatility of the seq2seq NLP paradigm, its application to source code generation presents unique challenges, such as syntactic validity and global semantic coherence. To address these challenges, extensions like code-aware search strategies [14] and syntax-aware edit-decoders [20] were proposed. Manipulating abstract syntax trees rather than raw code lines has shown promise in better capturing syntax [21]. Alamir et. al showed this by implementing transition-based parsing on ASTs [22] which went on to become a tool [23].
More recent works have used pre-trained Language Models of Code (CodeLMs) such as Liu et. al who used CodeBERT, CodeGPT, CodeT5, amongst others, for code updates. Their findings reveal that CodeLMs, while performing well without temporal constraints, struggle in realistic time-sensitive scenarios, particularly with generalizing to new projects, handling larger methods, and complex updates, often generating null updates [24]. One solution, PCREQ is a 6-component approach evaluated against popular LLMs such as GPT-4o and DeepSeek V3/R1 and outperforms by 18%-20% [25]. Nevertheless, automated program repair tasks have seen success with the use of not only LLMs, but multi-agent LLM systems when testing on the SWE-Bench dataset [26].
As such, in this paper, we address the challenge of code updates by conceptualizing it as a sequence of bug-fixing tasks. Our approach involves a multi-agent LLM system that systematically utilizes code summaries to identify necessary updates, applies them, and iteratively resolves any issues that arise until the code is successfully built and all unit tests pass. This framework also automatically identifies when to hand over to a human for manual intervention.
III Methodology
Our approach (depicted in Fig. 1) makes use of a novel change localisation mechanism called Meta-RAG. We introduced Meta-RAG as a means for efficient and accurate code editing mechanism for large code bases [27]. Meta-RAG works with condensing the codebase by summarising it into a structured natural language replica. This reduces the codebase size in tokens by a large extent (79.9% for SWE-bench Lite dataset [27]) which acts as a preprocessing step to help with information retrieval for change localisation in the codebase by reducing number of tokens to perform this task, as well as providing a more coherent large picture of the codebase. The main process, which is a customised ALMAS framework [9], is following a loop of making changes and resolving issues until the project builds and all tests are passing; the same process is followed by developers to do an upgrade.
III-A Preprocessing
Summary generation is performed as a preprocessing step. Summary agent takes in the repository and builds code summaries and stores them in a data structure that aligns with the code Abstract Syntax Tree (AST). Details of this process could be find in Tawosi et al. [27]. Briefly: LLM agent is asked to summarise each code file into a properly indented file, where each code unit (function or class) gets a singe line summary of its responsibility. Each file also gets a single line summary at file level which describes responsibility or purpure of the code file as a code unit. These LLM generated summaries are matched to the structure of the AST pasted from the code file to ensure correct mapping between the code units in the code file and the summary units in the summary file. These summaries are then traversed to find the relevant pieces of code for edit, thus using RAG on metadata instead of on the code itself. This summarization is done on the entire codebase as a one-off step, and only relevant parts are updated going forward as the codebase is updated.
III-B Main process
The main process (depicted in Fig. 1 inside the gray box) makes use of a number of agents and knowledge sources to complete the upgrade. We describe the agents and their responsibilities before diving into the process workflow.
III-B1 Control Agent
The control agent is another part of the Meta-RAG framework. Given a task, it is responsible for finding the code units “to read” that are useful as a context for a coding agent to perform the task, and the code units that are required to be modified (existing) or authored (new), i.e., “to write”. The context-to-read is to motivate the code agent to reuse the code from the repo and integrate the changes in a natural way to the code base. The control agent uses summaries and also a description of the current task to create instructions for code retrieval. In our solution, the control agent does this by checking file-level summaries first (one line summary per code file in the codebase) to select the relevant files. Then, checking the full summary of the selected files to code units that will form the context. Having the selected code summary units, the counterpart code units can then be fetched from the code files to put together a context for the code agent.
III-B2 Code Agent
The code agent could be any generic autonomous coding agent, in this paper we use GPT-4o. It receives a task and code units as context (to read, and to modify/write) and implements a change. The context is minimised in length by control agent, instead of multiple code search operations commonly used by other code editing agents. This agent implements the code directly into the code files and notifies the summary agent to update the summaries for the modified code, if needed. It is essential to keep the code summaries up-to-date with the codebase.
III-B3 Process Workflow
The upgrade process begins when the user initiates it. The target version can be specified by the user or automatically retrieved from the artifactory to update to the latest available version. The control agent accesses the code base and generates a structure for every pom and yml file in the project, as these files are not included in the summary base (which only summarizes code files). Additionally, the agent consults a migration guide for the target version, which provides general instructions for the upgrade. If a migration guide is unavailable, the control agent will rely on its internal knowledge to perform the upgrade. The migration guide may contain organization-specific changes necessary for the upgrade that are not publicly available, and which the LLM-based agent would not be aware of.
Using the available information, the control agent is tasked with producing initial upgrade instructions, deciding which files need modification and specifying the changes to be applied. The code agent then takes these instructions and implements the changes, which typically involve updating dependencies and altering configurations. After implementing the changes from the initial upgrade instructions, the project is compiled. If the compilation is successful, unit tests are executed. If all tests pass, the upgrade process is deemed successful. However, if the compilation fails or any unit tests fail, the error log is sent to the control agent for resolution. This process resembles Automated Program Repair (APR), where agents are tasked with editing the codebase to fix issues. The cycle of applying changes, compiling the code, running tests, and reviewing the output log continues until the build and test run is successful. The code agent indicates its inability to resolve the issue or if the same error occurs more than $n$ times (with $n=3$), the run stops. These conditions are implemented to prevent an infinite loop that could exhaust resources. This framework also supports collaboration, allowing AI agents to hand over the process to a human developer with a summary of the steps taken. If the developer resolves the issue and determines the process is incomplete, they can re-initiate it and return control to the AI agent to continue.
IV Upgrades Use Case
We evaluate our framework by applying it to upgrade a Java Moneta-based framework for three synthetic repositories. This section describes these repositories and the motivation behind evaluation with synthetic repositories. We also discuss the results of our framework in terms of similarity to the developer upgrade (i.e. Gold), and compare with an open-source agentic software development solution; OpenHands [28].
Figure 2: Illustration of synthetic code repositories used to evaluate the multi-agent LLM framework
TABLE I: Comparative results of the multi-agent LLM framework (LADU) against benchmark, showing the number of files, common files with Gold (G), lines added or removed, and common lines with (G) for each version upgrade scenario.
Approach Start Target #Files #File (G) #Common Files #Lines A:R #Lines A:R (G) #Common Lines
OpenHands + Claude 3.7 Sonnet 3.1.12 3.2.11 8 7 7 33:78 38:62 14:45
OpenHands + Claude 3.7 Sonnet 3.2.11 3.3.9 7 9 5 2:128 50:37 2:22
OpenHands + Claude 3.7 Sonnet 3.3.9 3.4.2 7 7 6 6:132 23:31 4:9
LADU + GPT-4o 3.1.12 3.2.11 8 7 7 37:67 38:62 13:48
LADU + GPT-4o 3.2.11 3.3.9 3 9 3 4:7 50:37 2:5
LADU + GPT-4o 3.3.9 3.4.2 3 7 3 10:9 23:31 4:5
TABLE II: Execution results of the multi-agent LLM framework (LADU) against benchmark.
Metric OpenHands LADU
3.1 3.2 3.3 3.1 3.2 3.3
Number of steps 106 86 72 18 16 4
Execution runtime [s] 1020.0 720.0 450.0 473.6 424.0 277.4
Total Tokens 1,514,456 822,781 594,366 78,421 54,812 14,387
Cost [\$] 4.66 2.38 1.83 0.62 0.45 0.11
IV-A Synthetic Dataset
The dependency to be upgraded is a cloud native event-sourced microservice framework for Java, built upon Spring Boot, and packaged into a dependency. Thousands of code repositories in industry use this framework. However, when an upgrade is made, not all required code updates are applied to all code repositories. As such, in order to test that our framework is able to make all updates covered in a particular upgrade, it would require an infrastructure setup that scales to hundreds of code repositories. In order to overcome this, we have created three synthetic code repositories that cover all changes required for an upgrade (Fig. 2).
The synthetic repositories are upgraded manually to act as a ”Gold” standard by which we compare our solution. As the synthetic repositories represent those we see in a realistic industrial setting, they include many yaml configuration files, pom files, in addition to the Java source code files.
IV-B Results
The framework is evaluated using the aforementioned three synthetic code repositories against the ”Gold” (G) standard, each representing different upgrade scenarios for the framework. The repositories are designed to test the framework’s ability to apply all necessary updates for specific version transitions: from versions 3.1 to 3.2, 3.2 to 3.3, and 3.3 to 3.4. We select a baseline (OpenHands) to benchmark against our LLM Agents for Dependency Upgrades (LADU).
Table I compares the framework’s performance with OpenHands + Claude 3.7 Sonnet and LADU + GPT-4o approaches. Key metrics include the number of files updated by each approach, the number of files in common with (G), the number of lines added or removed (A:R) by each approach, and common lines between the framework’s output and (G). The common file and line metrics are calculated as an exact match between the patch diffs. This is because LADU is run until unit tests pass, but the tests may not be comprehensive and do not cover deployment-related configurations in yaml files.
Table II presents a comparative analysis of the multi-agent LLM framework (LADU) against the OpenHands benchmark with regards to the execution metrics in order to assess efficiency. Number of steps, execution runtime in seconds, total tokens and cost in USD are reported.
IV-C Discussion
The results indicate that the proposed framework, while not perfectly matching the ”gold” standard, demonstrates a significant ability to automate dependency upgrades with a reasonable degree of accuracy. The number of common files and lines suggests that the framework can effectively identify and implement necessary changes, although there is room for improvement in aligning with manual upgrades.
The first version upgrade was presented with the most comprehensive migration guide, which is reflected in the performance of both approaches in Table I. LADU is able to achieve comparable performance to the baseline whilst using GPT-4o and comprising of a much simpler codebase and structure. More specifically, with regards to recall, our framework is able to make more correct line changes with regards to the addition of relevant lines, but does not remove as many correct ones as OpenHands.
Nonetheless, when looking at precision, LADU overall attempts fewer changes and thus maintains a higher precision, this is important in a realistic setting in order to avoid unwanted code changes in the repository during an upgrade. For instance, in upgrade 3.2 to 3.3, OpenHands removes 128 lines in the codebase, 22 of which are correct (17.2% precision), whilst LADU removes 7 lines, 5 of which were correct (71.4% precision). Thus our framework is able to have a higher confidence in making changes.
Table II results indicate that LADU maintains a significantly reduced number of steps required for the upgrade process, with only 18 steps needed for version 3.1 compared to 106 steps for OpenHands. Consequently, LADU also demonstrates a low execution runtime. Additionally, the framework uses fewer tokens, particularly in version 3.3, where it utilizes 14,387 tokens compared to OpenHands’ 594,366 tokens, due to the Meta-RAG mechanism that condenses the codebase. It is worth noting that the number of tokens used for summary generation are also included in LADU’s token count. Thus our approach is able to efficiently achieve comparable results to state-of-the-art benchmarks.
V Conclusion
This paper presents a novel framework consisting of multiple agents, including a Summary Agent, Control Agent, and Code Agent, which work together to update outdated library usages in codebases, while ensuring compatibility with new versions. The framework employs a change localization mechanism (Meta-RAG), which condenses the codebase, facilitating efficient information retrieval. The system was evaluated using synthetic code repositories, demonstrating its ability to automate upgrades with precision and efficiency compared to existing solutions. In the future, the evaluation can be expanded to include real-world codebases, and robust unit tests, to provide a more comprehensive assessment of the framework’s effectiveness in industrial settings. Furthermore, integrating more advanced LLMs and exploring hybrid approaches might also improve performance. Overall, the framework represents a promising advancement in automated software maintenance, offering a scalable solution for managing Java dependency upgrades in large codebases.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co and its affiliates (“JP Morgan”), and is not a product of the Research Department of JP Morgan. JP Morgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.